# The Indian Pulsar Timing Array Data Release 2: II. Customised Single-Pulsar Noise Analysis and Noise Budget

### v1

@@ 6 Discussion
The customised SPNA performed on the InPTA DR1 (hereafter, DR1-NA) dataset (Tarafdar+2022) was reported in Srivastava2023 for the sample of 14 pulsars over a time baseline of about $3.5$ years. The details of the adopted procedures are described in Srivastava2023, however there are a few differences from the present work that should be highlighted. The DR1-NA adopted a sequential model selection methodology, wherein, first the optimal white noise model was selected without any red noise model and then the optimal red noise model was selected using Bayes factors, including the white noise model obtained in the first step and keeping the Fourier basis fixed at $N_{\text{harm}}=N_{\text{max}}$. The optimal Fourier harmonics selection was done for a pre-defined grid with $k=2,5,8,12$ such that $f_{k}=k\times T_{\text{span}}^{\text{yr}}/T_{\text{span}}$, where $T_{\text{span}}^{\text{yr}}$ is the time span in years, which was also done in a sequential manner for different red noise processes. The adopted choices for white noise models did not include ECORR and red noise modeling predominantly included the scattering noise model (see §3.2.4). This differs significantly from the current work wherein a more generalised framework of optimal model and Fourier harmonics selection (see Section 3.3 for details) is incorporated along with the inclusion of ECORR in white noise modeling and FCN for additional chromatic modeling in place of the scattering model. Therefore, in this section, we restrict ourselves to a rather qualitative discussion of the two noise budgets.
@@
Broadly, the results of DR1-NA and the present work seem to be consistent with slight improvements – tighter constraints on ARN for PSRs J0613-0200 and J1643-1224, emergence of ARN in place of DMN for PSR J1012++5307 with better constraints on the spectral index that may have mimicked the DMN process in DR1-NA and steeper ARN spectral index for PSR J1939++2134 – owing to the extended time baseline of $\sim$$7.2\,\text{yr}$ in DR2 compared to $\sim$$3.5\,\text{yr}$ in DR1. In the present work, however, we do not find significance for ARN in PSRs J2124-3358 (having DMN) and J2145-0750 (having DMN++FCN), which also show significant non-Gaussianity levels (see Table 3). Such discrepancies may arise due to spectral contamination from unmodeled white noise in case of these pulsars in DR1-NA (Iraci+2024). However, a more rigorous analysis is required to gather a better understanding of the differences in the results, which we defer to a future work.
@@
One particular case to look at is that of PSR J1909-3744. In DR1-NA, a detailed analysis was presented to rule out the scattering model on physical grounds (see Section 5 in Srivastava2023). In the present work, however, we see a strong preference to the FCN model that seems to indicate otherwise. In order to assess this, we went back to the SW-cut dataset (see Section 5) and selected the ARN++DMN red noise model for the subsequent generation of post-noise residuals. We observed that the residuals had significant outliers near the solar-conjunction threshold epochs, which indicate residual solar-wind signatures that are, otherwise, modeled by the FCN model. This is not unexpected since the applied solar elongation threshold need not take care of all the solar-wind effects on the ToAs (Tiburzi+2019; Tiburzi2021). Therefore, the presence of FCN in the present work may not be related to an actual chromatic process in the ToAs, rather the impact of unmodeled solar-wind effects. This is also in agreement with the physical aspects detailed in Srivastava2023.
@@ 7 Summary
In this work, we presented the results of customised single-pulsar noise analysis on the pulsar timing residuals that were released as part of the second data release of the InPTA consortium (Rana+2025). We modeled various noises in the dataset in terms of white and red noise processes owing to their temporal characteristics. The white noise model uses three parameters, namely EFAC, EQUAD and ECORR, for different backend-receiver configurations, wherein the first one accounts for radiometer noise and the latter two take care of pulse jitter, all of which modify the final ToA uncertainties. The modeled red noise processes include achromatic red noise (ARN; radio-frequency independent but correlated across time), dispersion measure noise (DMN; varying with radio-frequency as $\nu^{-2}$ and correlated across time) and free chromatic noise (FCN; varying radio-frequency dependence and correlated across time), each of which is modeled as a stationary Gaussian process in the Fourier-domain with a power-law kernel. We model the solar-wind noise with a deterministic process with the average free-electron density $n_{\text{earth}}$ and its first time-derivative $\dot{n}_{\text{earth}}$ terms as free parameters. We performed model selection for multiple combinations of white noise and red noise processes using nested sampling (Skilling2004) to arrive at the optimal noise models that are summarised in Table 2. Fourier harmonics selection was performed with a dropout power-law implementation which uses a single hyper-parameter to effectively turn the noise process on/off depending on the Fourier basis cutoff, ultimately giving the suitable number of Fourier harmonics for the red noise process. Finally, we performed parameter estimation on the optimal noise model and Fourier harmonics to infer the optimal parameter estimates for different noise processes using MCMC sampling (Metropolis+1953; Hastings1970). The optimal harmonics and parameter estimation results are summarised in Table 5. We generated the noise-subtracted (post-noise) residuals by incorporating the effects of modeled red noise processes and scaling the ToA uncertainties with the estimated white noise parameters and these were subjected to the Anderson-Darling test for Gaussianity to assess the efficacy of the optimal noise models.
@@
We found that of the total of 27 pulsars, 15 have fully Gaussian, 8 have nearly Gaussian ($\text{p-value}<0.05$ and A-D statistic smaller than 2.5) while 4 have non-Gaussian ($\text{p-value}<10^{-4}$ and A-D statistic larger than 3.9) post-noise residuals (see Table 3 for details). We examined the potential causes that may lead to shortcomings in the current noise modeling landscape and coming out as non-Gaussianity in the noise-subtracted residuals. One prominent cause is the unmodeled solar-wind effect. In order to showcase this, we performed a separate analysis on pulsars that have strong solar-wind effects in the DMs and ToAs in the InPTA DR2, which included 4 pulsars with residual non-Gaussianity, and obtained sub-datasets with different ToA cut thresholds determined by a judicious trade-off between removal of solar-wind related artifacts in the DM time-series and timing residuals, at the same time, keeping the loss of total ToAs to less than 20%. We used the tension metrics approach from Raveri+2021 to quantify the parameter differences. We recovered significant parameter deviations between the two datasets with PSR J1022++1001 reaching as high as $3\sigma$ in tension for the DM noise process (see Table 4 for details). These results show that careful solar-wind modeling is crucial to obtain precise noise budget of a PTA, and the subsequent characterisation of the GWB, which itself manifests as a spatially correlated common red noise process and can be easily masked if other processes are not properly modeled. Furthermore, residual solar-wind effects can be mis-modeled by the FCN process in pulsars which do not show any observed chromatic effects in their pulse profiles originating from physical processes. We showcased this scenario for PSR J1909-3744 which strongly preferred the FCN model, despite the absence of any observed pulse broadening, as discussed in Srivastava2023. Therefore, careful modeling and its subsequent characterisation is crucial for assessing the overall PTA noise budget.
@@ 8 Future Work
The analysis methodology presented in this work, though thorough, need not be exhaustive. As discussed in Section 6, the existing non-Gaussianity levels for many pulsars point towards unmodeled effects in the timing residuals. We are currently looking into modeling white noise in a more efficient way to minimise leakages. A GP-based solar-wind noise model is being investigated to take care of the stochastic part of this process along with the deterministic model currently being used. We plan to examine the specific highly non-Gaussian pulsars (see Section 6) in more detail to identify and address the potential sources that may give rise to such behavior. There is also a need to model the chromatic noise process in a better way so as to prevent any contamination of ARN that may cause potential bias in the estimates of the GWB. All of these are currently being investigated as part of a future work (Dwivedi et al., in preparation). These attempts at rigorous modeling of single-pulsar noise sources are crucial in preventing false-positives in GWB detections and its subsequent characterisation, along with unveiling the presence of any other potential astrophysical signals in the dataset.
@@ Acknowledgements
InPTA acknowledges the support of the uGMRT staff in resolving technical issues and telescope operators for the observations. The uGMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research, India. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM-Ganga’ at the Indian Institute of Technology Roorkee, as well as ‘PARAM-Seva’ at IIT Hyderabad, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. CD acknowledges the Param Vikram-1000 High Performance Computing Cluster of the Physical Research Laboratory (PRL), a unit of the Department of Space, Government of India, for performing the intensive computations. BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886 and 1002/2/2023/RRC/R&D-II/14369). SD is supported by ANRF MTR/2023/000384. KT is partially supported by JSPS KAKENHI Grant Numbers 20H00180, 21H01130, 21H04467, 24H01813, and 25K21670 and Bilateral Joint Research Projects of JSPS 120237710.380. PR acknowledges the financial assistance of the South African Radio Astronomy Observatory (SARAO) towards this research (www.sarao.ac.za). PA, BCJ, HG and ASh acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Ganga’ at the Indian Institute of Technology Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. KR is supported by CSIR NET JRF. JS acknowledges funding from the South African Research Chairs Initiative of the Department of Science and Technology and the National Research Foundation of South Africa. HT is supported by DST INSPIRE Fellowship. AKP is supported by CSIR fellowship Grant number 09/0079(15784)/2022-EMR-I. ZZ is supported by the Prime Minister’s Research Fellows (PMRF) scheme, Ref. No. TF/PMRF-22-7307. KV is supported by CSIR JRF Fellowship. ASh is supported by UGC JRF fellowship. NDB is supported by DST-WISE Post Doctoral Fellowship (DST/WISE-PDF/PM-17/2024(G). RK is supported by JSPS KAKENHI Grant Number 24K17051. DD acknowledges the usage of computational resources at the Institute of Mathematical Science’s High Performance Computing facility (hpc.imsc.res.in) [Kamet]
@@
Python (python), Enterprise (Ellis+2020; Johnson+2024), Enterprise Extensions (enterpriseext), Singularity (Singularity), Dynesty (Speagle2020), PTMCMCsampler (Ellis+2017), Tensiometer (Raveri2019), NumPy (NumPy), AstroPy (astropy), SciPy (SciPy), Matplotlib (Matplotlib), Corner (corner), StatsModels (statsmodels)

### v2

@@ 6 Discussion
The customised SPNA performed on the InPTA DR1 (hereafter, DR1-NA) dataset (Tarafdar et al., 2022) was reported in Srivastava et al. (2023) for the sample of 14 pulsars over a time baseline of about $3.5$ years. The details of the adopted procedures are described in Srivastava et al. (2023), however there are a few differences from the present work that should be highlighted. The DR1-NA adopted a sequential model selection methodology, wherein, first the optimal white noise model was selected without any red noise model and then the optimal red noise model was selected using Bayes factors, including the white noise model obtained in the first step and keeping the Fourier basis fixed at $N_{\text{harm}}=N_{\text{max}}$. The optimal Fourier harmonics selection was done for a pre-defined grid with $k=2,5,8,12$ such that $f_{k}=k\times T_{\text{span}}^{\text{yr}}/T_{\text{span}}$, where $T_{\text{span}}^{\text{yr}}$ is the time span in years, which was also done in a sequential manner for different red noise processes. The adopted choices for white noise models did not include ECORR and red noise modeling predominantly included the scattering noise model (see §3.2.4). This differs significantly from the current work wherein a more generalized framework of optimal model and Fourier harmonics selection (see Section 3.3 for details) is incorporated along with the inclusion of ECORR in white noise modeling and FCN for additional chromatic modeling in place of the scattering model. Therefore, in this section, we restrict ourselves to a rather qualitative discussion of the two noise budgets.
@@
Broadly, the results of DR1-NA and the present work seem to be consistent with slight improvements – tighter constraints on ARN for PSRs J0613-0200 and J1643-1224, emergence of ARN in place of DMN for PSR J1012++5307 with better constraints on the spectral index that may have mimicked the DMN process in DR1-NA and steeper ARN spectral index for PSR J1939++2134 – owing to the extended time baseline of $\sim$$7.2\,\text{yr}$ in DR2 compared to $\sim$$3.5\,\text{yr}$ in DR1. In the present work, however, we do not find significance for ARN in PSRs J2124-3358 (having DMN) and J2145-0750 (having DMN++FCN), which also show significant non-Gaussianity levels (see Table 3). Such discrepancies may arise due to spectral contamination from unmodeled white noise in case of these pulsars in DR1-NA (Iraci et al., 2024). However, a more rigorous analysis is required to gather a better understanding of the differences in the results, which we defer to a future work.
@@
One particular case to look at is that of PSR J1909-3744. In DR1-NA, a detailed analysis was presented to rule out the scattering model on physical grounds (see Section 5 in Srivastava et al. (2023)). In the present work, however, we see a strong preference for the FCN model that seems to indicate otherwise. In order to assess this, we went back to the SW-cut dataset (see Section 5) and selected the ARN++DMN red noise model for the subsequent generation of post-noise residuals. We observed that the residuals had significant outliers near the solar-conjunction threshold epochs, which indicate residual solar-wind signatures that are, otherwise, modeled by the FCN model. This is not unexpected since the applied solar elongation threshold need not take care of all the solar-wind effects on the ToAs (Tiburzi et al., 2019, 2021). Therefore, the presence of FCN in the present work may not be related to an actual chromatic process in the ToAs, rather the impact of unmodeled solar-wind effects. This is also in agreement with the physical aspects detailed in Srivastava et al. (2023).
As can be seen from Figure 4, there seems to be an apparent bimodality in the distribution of amplitude and chromatic indices of the pulsars, where curiously PSRs J1643-1224 and J0437-4715 show a very small chromatic index and relatively large amplitude. This behavior is, however, not seen for ARN and DMN processes (see Figures 2 and 3). This peculiar diversity in the results of FCN modeling may be due to different sensitivities in Band 3 and Band 5 (see Table 5) and is currently being investigated as part of a future work.
@@ 7 Summary
In this work, we presented the results of customised single-pulsar noise analysis on the pulsar timing residuals that were released as part of the second data release of the InPTA consortium (Rana et al., 2025a). We modeled various noises in the dataset in terms of white and red noise processes owing to their temporal characteristics. The white noise model uses three parameters, namely EFAC, EQUAD and ECORR, for different backend-receiver configurations, wherein the first one accounts for radiometer noise and the latter two take care of pulse jitter, all of which modify the final ToA uncertainties. The modeled red noise processes include achromatic red noise (ARN; radio-frequency independent but correlated across time), dispersion measure noise (DMN; varying with radio-frequency as $\nu^{-2}$ and correlated across time) and free chromatic noise (FCN; varying radio-frequency dependence and correlated across time), each of which is modeled as a stationary Gaussian process in the Fourier-domain with a power-law kernel. We model the solar-wind noise with a deterministic process with the average free-electron density $n_{\text{earth}}$ and its first time-derivative $\dot{n}_{\text{earth}}$ terms as free parameters. We performed model selection for multiple combinations of white noise and red noise processes using nested sampling (Skilling, 2004) to arrive at the optimal noise models that are summarized in Table 2. Fourier harmonics selection was performed with a dropout power-law implementation which uses a single hyper-parameter to effectively turn the noise process on/off depending on the Fourier basis cutoff, ultimately giving the suitable number of Fourier harmonics for the red noise process. Finally, we performed parameter estimation on the optimal noise model and Fourier harmonics to infer the optimal parameter estimates for different noise processes using MCMC sampling (Metropolis et al., 1953; Hastings, 1970). The optimal harmonics and parameter estimation results are summarised in Table 5. We generated the noise-subtracted (post-noise) residuals by incorporating the effects of modeled red noise processes and scaling the ToA uncertainties with the estimated white noise parameters and these were subjected to the Anderson-Darling test for Gaussianity to assess the efficacy of the optimal noise models.
@@
We found that of the total of 27 pulsars, 15 have fully Gaussian, 8 have nearly Gaussian ($\text{p-value}<0.05$ and A-D statistic smaller than 2.5) while 4 have non-Gaussian ($\text{p-value}<10^{-4}$ and A-D statistic larger than 3.9) post-noise residuals (see Table 3 for details). We examined the potential causes that may lead to shortcomings in the current noise modeling landscape and coming out as non-Gaussianity in the noise-subtracted residuals. One prominent cause is the unmodeled solar-wind effect. In order to showcase this, we performed a separate analysis on pulsars that have strong solar-wind effects in the DMs and ToAs in the InPTA DR2, which included 4 pulsars with residual non-Gaussianity, and obtained sub-datasets with different ToA cut thresholds determined by a judicious trade-off between removal of solar-wind related artifacts in the DM time-series and timing residuals, at the same time, keeping the loss of total ToAs to less than 20%. We used the tension metrics approach from Raveri and Doux (2021) to quantify the parameter differences. We recovered significant parameter deviations between the two datasets with PSR J1022++1001 reaching as high as $3\sigma$ in tension for the DM noise process (see Table 4 for details). These results show that careful solar-wind modeling is crucial to obtain precise noise budget of a PTA, and the subsequent characterization of the GWB, which itself manifests as a spatially correlated common red noise process and can be easily masked if other processes are not properly modeled. Furthermore, residual solar-wind effects can be mis-modeled by the FCN process in pulsars which do not show any observed chromatic effects in their pulse profiles originating from physical processes. We showcased this scenario for PSR J1909-3744 which strongly preferred the FCN model, despite the absence of any observed pulse broadening, as discussed in Srivastava et al. (2023). Therefore, careful modeling and its subsequent characterization is crucial for assessing the overall PTA noise budget.
At this juncture, it is worthwhile to point out that while the analysis presented in this work is broadly consistent with what is done by other PTA consortia, it is still not straightforward to directly compare the noise budgets of different PTAs (even for the overlapping set of pulsars). The noise estimates are highly dependent on the sensitivities of individual PTA datasets along with the methodologies incorporated to model them. The DMX-based chromatic noise modeling approach adopted by NANOGrav (Agazie et al., 2023b), the inclusion of additional System and Band Noises incorporated by PPTA (Reardon et al., 2023b), the different normalization schemes adopted in EPTA for the DM process (EPTA Collaboration et al., 2023) and the overall varying functional form of incorporating white noise parameters, poses a great difficulty in the interpretation of inter-PTA noise budgets. Therefore, in this work, we do not present any such study and an investigation on the same grounds as done in Agazie et al. (2024) is essential to arrive at such comparisons.
@@ 8 Future Work
The analysis methodology presented in this work, though thorough, need not be exhaustive. As discussed in Section 6, the existing non-Gaussianity levels for many pulsars point towards unmodeled effects in the timing residuals. We are currently looking into modeling white noise in a more efficient way to minimize leakages. A GP-based solar-wind noise model is being investigated to take care of the stochastic part of this process along with the deterministic model currently being used. We plan to examine the specific highly non-Gaussian pulsars (see Section 6) in more detail to identify and address the potential sources that may give rise to such behavior. There is also a need to model the chromatic noise process in a better way so as to prevent any contamination of ARN that may cause potential bias in the estimates of the GWB. All of these are currently being investigated as part of a future work (Dwivedi et al., in preparation). These attempts at rigorous modeling of single-pulsar noise sources are crucial in preventing false-positives in GWB detections and its subsequent characterization, along with unveiling the presence of any other potential astrophysical signals in the dataset.
@@ Acknowledgements
InPTA acknowledges the support of the uGMRT staff in resolving technical issues and telescope operators for the observations. The uGMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research, India. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM-Ganga’ at the Indian Institute of Technology Roorkee, as well as ‘PARAM-Seva’ at IIT Hyderabad, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. CD acknowledges the Param Vikram-1000 High Performance Computing Cluster of the Physical Research Laboratory (PRL), a unit of the Department of Space, Government of India, for performing the intensive computations. The work of CD at the Physical Research Laboratory (PRL) was supported by the Department of Space, Government of India. BCJ acknowledges the support from Raja Ramanna Chair fellowship of the Department of Atomic Energy, Government of India (RRC – Track I Grant 3/3401 Atomic Energy Research 00 004 Research and Development 27 02 31 1002//2/2023/RRC/R&D-II/13886 and 1002/2/2023/RRC/R&D-II/14369). SD is supported by ANRF MTR/2023/000384. KT is partially supported by JSPS KAKENHI Grant Numbers 20H00180, 21H01130, 21H04467, 24H01813, and 25K21670 and Bilateral Joint Research Projects of JSPS 120237710.380. PR acknowledges the financial assistance of the South African Radio Astronomy Observatory (SARAO) towards this research (www.sarao.ac.za). PA, BCJ, HG and ASh acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Ganga’ at the Indian Institute of Technology Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. KR is supported by CSIR NET JRF. JS acknowledges funding from the South African Research Chairs Initiative of the Department of Science and Technology and the National Research Foundation of South Africa. HT is supported by DST INSPIRE Fellowship. AKP is supported by CSIR fellowship Grant number 09/0079(15784)/2022-EMR-I. ZZ is supported by the Prime Minister’s Research Fellows (PMRF) scheme, Ref. No. TF/PMRF-22-7307. KV is supported by CSIR JRF Fellowship. ASh is supported by UGC JRF fellowship. NDB is supported by DST-WISE Post Doctoral Fellowship (DST/WISE-PDF/PM-17/2024(G). RK is supported by JSPS KAKENHI Grant Number 24K17051. DD acknowledges the usage of computational resources at the Institute of Mathematical Science’s High Performance Computing facility (Kamet). Finally, we would like to thank the referee for many useful and constructive comments on the manuscript.
@@
Python (Van Rossum and Drake, 2009), Enterprise (Ellis et al., 2020; Johnson et al., 2024), Enterprise Extensions (Taylor et al., 2021), Singularity (Kurtzer et al., 2017), Dynesty (Speagle, 2020), PTMCMCsampler (Ellis and van Haasteren, 2017), Tensiometer (Raveri and Hu, 2019), NumPy (Harris et al., 2020), AstroPy (Astropy Collaboration et al., 2022), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), Corner (Foreman-Mackey, 2016), StatsModels (Seabold and Perktold, 2010)